# Quasirandom and quasisimple groups

### v1

@@ 4.2. Proof of Theorem B
From the data collected in the Appendix, we note that, among the finite simple groups, $J_{1}$ and $O^{\prime}N$ are the only ones whose quasirandom degree exceeds $\frac{1}{3}$. We claim that no other group has this property. Looking for a contradiction, suppose that $G$ is of minimal cardinality among the finite groups satisfying $\bm{\varepsilon}(G)\geq\frac{1}{3}$ that are not isomorphic to either $J_{1}$ or $O^{\prime}N$. Observe that $G$ is not simple, and let $N\lhd G$ be a nontrivial proper normal subgroup. We have $\bm{\varepsilon}(G/N)\geq\frac{1}{3}$, and $G/N$ is isomorphic to either $J_{1}$ or $O^{\prime}N$. In the former case, from (2.1), we have
@@
We are ready to prove Theorem B. This proof has a number of cases, the most difficult, which we treat in the end, being the following. We have an abelian minimal normal subgroup $N$ of $G$ such that $G/N$ is quasisimple, and the action of $G/N$ on $N$ by conjugation provides a projective modular representation of the simple quotient of $G/N$, say $L$. The key idea is that the (complex, linear) quasirandomness of $L$ controls also the degrees of the modular projective representations. This gives a lower bound on $|N|$, which we put against the upper bound given by (2.1). We start with the details.
@@
where $N$ is the unique minimal normal subgroup of $G$, and $M$ is the unique maximal normal subgroup of $G$. It follows that $G/M$ is isomorphic to a simple group $L$ satisfying $\bm{\varepsilon}(L)>\frac{1}{5}$, and that $N\cong T^{n}$ is the direct product of isomorphic simple groups. We distinguish two cases: either $G/N$ is quasisimple or not. Let us start from the latter. $G/N$ is not quasisimple. We have $\bm{\varepsilon}(G/N)>\frac{1}{5}$, and, by the minimality of $G$, we are in one of the two cases:
@@ 2. $(b)$
This is impossible. It remains the case where $N$ is nonabelian, say $N=T^{n}$ where $T$ is a nonabelian finite simple group. We argue in a similar way as in Theorem A: we don’t have $C=\mathbf{C}_{G}(N)\subseteq N$ as before, but the same argument works because $C\subseteq M$. We have
(4.2) $|G|\leq\mathbf{D}_{0}(\mathrm{Sp}_{4}(q))^{5}.$
Considering the embedding of $G/C$ into $\mathrm{Aut}(T^{n})$, and arguing as before, we obtain $\mathbf{D}_{0}(\mathrm{Sp}_{4}(q))\leq n$. Since $|N|\geq 60^{n}$, and $60^{x}>x^{5}$ for every $x\geq 1$, this contradicts (4.2). $G/N$ is quasisimple. Let $C=\mathbf{C}_{G}(N)$. We distinguish three cases. Suppose that $C=G$. Here $N$ is central and $G$ is quasisimple by Lemma 2.10, a contradiction. Suppose that $C=1$. Here $N=T^{n}$ is the product of nonabelian finite simple groups, and $G$ is embedded into $\mathrm{Aut}(T^{n})$. Let $G/M=L$. Arguing as in the proof of Theorem A, we obtain
@@
This is impossible for any value of $\mathbf{D}_{0}(L)$. Suppose that $N\subseteq C\subseteq M$. This case will take up the rest of the proof. First, Lemma 2.7 provides
@@
Moreover, $N$ is elementary abelian, and the action of $G/N$ on $N$ by conjugation provides a projective modular representation of $L$ (recall that $G/N$ is quasisimple). As already explained, the key idea is that the quasirandomness of $G/N$ gives a lower bound on $|N|$, which we put against the upper bound given by (2.1). We observe that $\mathrm{Alt}(5)$ and $\mathrm{Alt}(6)$ are $\frac{1}{4}$-quasirandom, but if one of them appears as $L$, then
@@ 1. $(i)$
    $L$ is a simple group of Lie type defined over a field of characteristic $p$, and $G/N$ induces a projective modular representation of $L$ in cross-characteristic;
@@ 2. $(ii)$
    $L$ is a simple group of Lie type defined over a field of characteristic $p$, and $G/N$ induces a projective modular representation of $L$ in natural characteristic;
@@ 3. $(iii)$
(4.3) $|N|\leq|M|\leq\frac{\mathbf{D}_{0}(L)^{5}}{|L|}<\mathbf{D}_{0}(L)^{2}.$
From Remark 2.14, the definition of $\mathbf{R}_{r}(L)$ with $r\neq p$ a prime, and (4.3),
@@
This case can be handled computationally (see Remark 4.8). Let $L$ be as described in $(ii)$. We first deal with $L\cong\mathrm{Sp}_{4}(q)$ with $q$ even, $q\geq 4$. Since the Schur multiplier is trivial, we have $N=M$. By (2.1), we find
@@
In all these four cases, the chain of inequalities is impossible. Let $L$ be as described in $(iii)$. Since $\bm{\varepsilon}(L)>\frac{1}{4}$, the possibilities for $L$ are
@@
In the following tables, we collect the values $\mathbf{D}_{0}(L)$ and $\bm{\varepsilon}(L)$ when $L$ is a finite simple group of Lie type or a sporadic simple group. While gathering these data, we stumbled into a little gap in the literature: as far as we know, the minimal degree of the complex irreducible representations for classical simple groups have never been computed and recorded in a single place. Let $L$ be a classical group of Lie type. In [28], Tiep and Zalesskii computed the minimal dimension of a projective complex irreducible representation of $L$ (that is, a complex irreducible representation of $\widetilde{L}$). In Table 2, we record $\mathbf{D}_{0}(L)$ for the classical groups of Lie type.
These data were virtually known, in the sense that all the complex representation of minimal dimension were already described as projective representations whose restrictions on the center are trivial. First, we notice that
$\mathrm{PSL}_{3}(2),\>\mathrm{PSL}_{3}(4),\>\mathrm{PSL}_{4}(2),\>\mathrm{PSL}_{4}(3),\>\mathrm{PSU}_{4}(2),\>\mathrm{PSU}_{4}(4),\>\Omega_{7}(3),\>\Omega^{+}_{8}(2)$
can be dealt with a calculator, while the character table of $\mathrm{PSL}_{2}(q)$ is well known. Now, note that unipotent representations of quasisimple groups are trivial on their center. Hence, using [28, Theorems 3.1, 4.1, 5.2, 7.6] and [17, Theorem 2.5], we obtain that for $\mathrm{PSL}_{d}(q)$, $\mathrm{PSU}_{d}(q)$ with $d$ odd, $\mathrm{Sp}_{2d}(q)$ with $q$ even, and $\mathrm{P}\Omega_{2d}^{\pm}(q)$, $\mathbf{D}_{0}(L)$ coincide with the minimal dimension of a projective unipotent representation. We are left to deal with $\mathrm{PSU}_{d}(q)$ with even rank, $\mathrm{PSp}_{2d}(q)$ in odd characteristic, and $\Omega_{2d+1}(q)$. We start with $\mathrm{PSU}_{d}(q)$ with $d$ even. By [28, Theorem 4.1], the second minimal dimension of a complex representation $\rho$ of $\mathrm{SU}_{d}(q)$ is
$\dim(\rho)=\mathbf{D}_{0}(\mathrm{SU}_{d}(q))+1=\frac{q^{d}+q}{q+1}.$
Further, such a representation is unipotent. Hence, $\mathbf{D}_{0}(\mathrm{PSU}_{d}(q))=\dim(\rho)$. Let us tackle $\mathrm{PSp}_{2d}(q)$ with $q$ odd. In view of [28, Theorem 5.2], the unitary representations $\rho_{+}$ and $\rho_{-}$ with smallest dimension of $\mathrm{Sp}_{2d}(q)$ are such that
$\dim(\rho_{-})=\frac{q^{d}-1}{2},\quad\hbox{and}\quad\dim(\rho_{+})=\frac{q^{d}+1}{2}.$
Observe that, as $\dim(\rho_{+})=\dim(\rho_{-})+1$, one of these representations has even degree, while the other has odd degree. The latter is trivial on the center, therefore it projects to an irreducible representation of $\mathrm{PSp}_{2d}(q)$ of minimal degree. Finally, we must consider $\Omega_{2d+1}(q)$. From [17, Theorem 2.2], the two unitary representations of $\mathrm{Spin}_{2d+1}(q)$ with minimal degrees are such that
$\dim(\rho_{1})=\frac{q^{2d}-1}{q^{2}-1},\quad\hbox{and}\quad\dim(\rho_{2})=\frac{(q^{d}-1)(q^{d}-q)}{2(q+1)}.$
Observe that $\rho_{2}$ is unipotent, while $\rho_{1}$ is not. For $q=3$, $\dim(\rho_{1})\leq\dim(\rho_{2})$, so we find that $\mathbf{D}_{0}(\Omega_{2d+1}(3))=\dim(\rho_{2})$. For all the other powers of odd primes, the inequality is reversed. As before, if $\dim(\rho_{1})$ is odd, $\rho_{1}$ is trivial on the center, thus $\mathbf{D}_{0}(\Omega_{2d+1}(q))=\dim(\rho_{1})$. Otherwise, we have $\mathbf{D}_{0}(\Omega_{2d+1}(q))=\dim(\rho_{2})$.
@@
$d\geq 4$ even, $(d,q)\neq(4,2)$ $\dfrac{q^{d}+q}{q+1}$
@@
$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{Sp}_{2d}(q)$ $d\geq 2$, $q$ even $\dfrac{(q^{d}-1)(q^{d}-q)}{2(q+1)}$
@@
Table 2. Minimal degree of a representation of simple classical groups.
In Tables 3, 4, and 5, we collect the asymptotic quasirandom degrees of the finite simple groups of Lie type, and the quasirandom degrees of the sporadic simple groups. These informations are collected from [7, 13, 14, 15, 17, 25, 27, 28].
$\dfrac{1^{1}}{1_{1_{1}}}$ $L$ Comments Asymptotic for $|L|$ Asymptotic for $\mathbf{D}_{0}(L)$ $\lim\limits_{q\rightarrow+\infty}\bm{\varepsilon}(L)$
$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{PSL}_{d}(q)$ $d\geq 2$ $q^{d^{2}-1}$ $q^{d-1}$ $\dfrac{1}{d+1}$
$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{PSU}_{d}(q)$ $d\geq 3$ $q^{d^{2}-1}$ $q^{d-1}$ $\dfrac{1}{d+1}$
$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{Sp}_{2d}(q)$ $d\geq 2$, $q$ even $q^{2d^{2}+d}$ $q^{2d-1}$ $\dfrac{2d-1}{d(2d+1)}$
$d\geq 2$, $q$ odd $q^{2d^{2}+d}$ $q^{d}$ $\dfrac{1}{2d+1}$
$\dfrac{1^{1}}{1_{1_{1}}}$ $\Omega_{2d+1}(q)$ $d\geq 3$, $q$ odd $q^{2d^{2}+d}$ $q^{2d-2}$ $\dfrac{2(d-1)}{d(2d+1)}$
$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{P}\Omega^{\pm}_{2d}(q)$ $d\geq 4$ $q^{2d^{2}-d}$ $q^{2d-3}$ $\dfrac{2d-3}{d(2d-1)}$
Table 3. Asymptotic quasirandom degree of simple classical groups of Lie type.
$\dfrac{1^{1}}{1_{1_{1}}}$ $L$ Comments Asymptotic for $|L|$ Asymptotic for $\mathbf{D}_{0}(L)$ $\lim\limits_{q\rightarrow+\infty}\bm{\varepsilon}(L)$
$\dfrac{1^{1}}{1_{1_{1}}}$ $E_{6}(q)$ $q^{78}$ $q^{11}$ $\dfrac{11}{78}$
$\dfrac{1^{1}}{1_{1_{1}}}$ $E_{7}(q)$ $q^{133}$ $q^{17}$ $\dfrac{17}{133}$
$\dfrac{c1^{1}}{1_{1_{1}}}$ $E_{8}(q)$ $q^{248}$ $q^{29}$ $\dfrac{29}{248}$
$\dfrac{1^{1}}{1_{1_{1}}}$ $F_{4}(q)$ $q$ even $q^{52}$ $q^{11}$ $\dfrac{11}{52}$
$\dfrac{1^{1}}{1_{1_{1}}}$ $F_{4}(q)$ $q$ odd $q^{52}$ $q^{8}$ $\dfrac{2}{13}$
@@
$\dfrac{1^{1}}{1_{1_{1}}}$ ${}^{2}E_{6}(q^{2})$ $q^{78}$ $q^{11}$ $\dfrac{11}{78}$
$\dfrac{1^{1}}{1_{1_{1}}}$ ${}^{3}D_{4}(q^{3})$ $q^{28}$ $q^{5}$ $\dfrac{5}{28}$
$\dfrac{1^{1}}{1_{1_{1}}}$ ${}^{2}F_{4}(q^{2})$ $q=\sqrt{2^{2m+1}}$ $q^{52}$ $q^{11}$ $\dfrac{11}{52}$
$\dfrac{1^{1}}{1_{1_{1}}}$ ${}^{2}B_{2}(q^{2})$ $q=\sqrt{2^{2m+1}}$ $q^{10}$ $q^{3}$ $\dfrac{3}{10}$
$\dfrac{1^{1}}{1_{1_{1}}}$ ${}^{2}G_{2}(q^{2})$ $q=\sqrt{3^{2m+1}}$ $q^{14}$ $q^{4}$ $\dfrac{2}{7}$
Table 4. Asymptotic quasirandom degree of simple exceptional groups of Lie type.

### v2

@@ 4.2. Proof of Theorem B
From the data collected in the Appendix, we note that, among the finite simple groups, $J_{1}$ and $O^{\prime}N$ are the only ones whose quasirandom degree is at least $\frac{1}{3}$. Suppose that $G$ is of minimal cardinality among the finite groups satisfying $\bm{\varepsilon}(G)\geq\frac{1}{3}$ that are not isomorphic to either $J_{1}$ or $O^{\prime}N$. Of course, there exists a nontrivial proper normal subgroup $N$ of $G$. We have $\bm{\varepsilon}(G/N)\geq\frac{1}{3}$, and $G/N$ is isomorphic to either $J_{1}$ or $O^{\prime}N$. In the former case, from (2.1), we have
@@
We are ready to prove Theorem B. This proof has a number of cases, the most difficult, which we treat in the end, being the following. We have an abelian minimal normal subgroup $N$ of $G$ such that $G/N$ is quasisimple, and the action of $G/N$ on $N$ by conjugation provides a projective modular representation of the simple quotient of $G/N$, say $L$. The key point is that the (complex, linear) quasirandomness of $L$ controls also the degrees of the modular projective representations. This gives a lower bound on $|N|$, which we put against the upper bound given by (2.1). We start with the details.
@@
where $N$ is the unique minimal normal subgroup of $G$, and $M$ is the unique maximal normal subgroup of $G$. It follows that $G/M$ is isomorphic to a simple group $L$ satisfying $\bm{\varepsilon}(L)>\frac{1}{5}$, and that $N\cong T^{n}$ is the direct product of isomorphic simple groups. We distinguish two cases: either $G/N$ is quasisimple or not. Let us start from the latter.
$G/N$ is not quasisimple. We have $\bm{\varepsilon}(G/N)>\frac{1}{5}$, and, by the minimality of $G$, we are in one of the two cases:
@@ 2. $(b)$
This is impossible. It remains the case where $N$ is nonabelian, say $N=T^{n}$ where $T$ is a nonabelian finite simple group. We argue in a similar way as in Theorem A: since $N$ is the unique minimal normal subgroup of $G$, we have $\mathbf{C}_{G}(N)=1$ and the same argument works. We have
(4.2) $|G|\>\leq\>\mathbf{D}_{0}(\mathrm{Sp}_{4}(q))^{5}.$
Considering the embedding of $G$ into $\mathrm{Aut}(T^{n})$, and arguing as before, we obtain $\mathbf{D}_{0}(\mathrm{Sp}_{4}(q))\leq n$. Since $|N|\geq 60^{n}$, and $60^{x}>x^{5}$ for every $x\geq 1$, this contradicts (4.2).
$G/N$ is quasisimple. Let $C=\mathbf{C}_{G}(N)$. We distinguish three cases.
Suppose that $C=G$. Here $N$ is central and $G$ is quasisimple by Lemma 2.10, a contradiction.
Suppose that $C=1$. Here $N=T^{n}$ is the product of nonabelian finite simple groups, and $G$ is embedded into $\mathrm{Aut}(T^{n})$. Let $G/M=L$. Arguing again as in the proof of Theorem A, we obtain
@@
This is impossible for any value of $\mathbf{D}_{0}(L)$.
Suppose that $N\subseteq C\subseteq M$. This case will take up the rest of the proof. First, $N$ is an elementary abelian noncentral subgroup, and Lemma 2.7 provides
@@
Moreover, the action of $G/N$ on $N$ by conjugation provides a projective modular representation of $L$ (recall that $G/N$ is quasisimple). As already explained, the key idea is that the quasirandomness of $G/N$ gives a lower bound on $|N|$, which we put against the upper bound given by (2.1). We observe that $\mathrm{Alt}(5)$ and $\mathrm{Alt}(6)$ are $\frac{1}{4}$-quasirandom, but if one among them appears as $L$, then
@@ 1. $(i)$
    $L$ is a simple group of Lie type over a field of characteristic $p$, and $G/N$ induces a projective modular representation of $L$ in cross-characteristic;
@@ 2. $(ii)$
    $L$ is a simple group of Lie type over a field of characteristic $p$, and $G/N$ induces a projective modular representation of $L$ in natural characteristic;
@@ 3. $(iii)$
$|N|\leq|M|\leq\frac{\mathbf{D}_{0}(L)^{5}}{|L|}<\mathbf{D}_{0}(L)^{2}.$
From Remark 2.14, and the definition of $\mathbf{R}_{r}(L)$ with $r\neq p$ a prime, we obtain
@@
This case can be handled computationally.
Let $L$ be as described in $(ii)$. We first deal with $L\cong\mathrm{Sp}_{4}(q)$ with $q$ even, $q\geq 4$. Since the Schur multiplier is trivial, we have $N=M$. By (2.1), we find
@@
In all these four cases, the chain of inequalities is impossible.
Let $L$ be as described in $(iii)$. Since $\bm{\varepsilon}(L)>\frac{1}{4}$, the possibilities for $L$ are
@@
In the following tables, we collect the values $\mathbf{D}_{0}(L)$ and $\bm{\varepsilon}(L)$ when $L$ is a finite simple group of Lie type or a sporadic simple group. For a group of Lie type $L_{d}(q)$, we fix $d$ and make $q$ go to infinity. The case where $L$ is of exceptional type is discussed in detail by Lübeck [17]. So let $L$ be a classical group of Lie type. In [27], Tiep and Zalesskii computed the minimal dimension of a projective complex irreducible representation of $L$, which we denote by $\mathbf{R}_{0}(L)$. See also [18]. In Table 2, we use these informations to record $\mathbf{D}_{0}(L)$ for the simple classical groups. We remark that $\mathbf{R}_{0}(L)<\mathbf{D}_{0}(L)$ holds in some cases, for example for some symplectic groups in odd characteristic.
@@
$d\geq 4$ even, $(d,q)\neq(4,2)$ $\dfrac{q^{d}-1}{q+1}\leq\mathbf{D}_{0}(L)\leq\dfrac{q^{d}+q}{q+1}$
@@
$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{PSp}_{2d}(q)$ $d\geq 2$, $q$ even $\dfrac{(q^{d}-1)(q^{d}-q)}{2(q+1)}$
@@
Table 2. $\mathbf{D}_{0}(L)$ for simple classical groups.
In Tables 3, 4, and 5, we collect the asymptotic quasirandom degrees of the finite simple groups of Lie type, and the quasirandom degrees of the sporadic simple groups [7]. We remark that, after $(\mathrm{PSL}_{2}(q))_{q\geq 4}$, which is asymptotically $\frac{1}{3}$-quasirandom, the most quasirandom families of finite simple groups are $(\mathrm{Sp}_{4}(q))_{q\,even}$ and the Suzuki groups $(^{2}B_{2}(q^{2}))_{q\geq 8}$, whose quasirandom degrees are asymptotically $\frac{3}{10}$.
$\dfrac{1^{1}}{1_{1_{1}}}$ $L$ Comments $\sim\log_{q}|L|$ $\sim\log_{q}\mathbf{D}_{0}(L)$ $\lim\limits_{q\rightarrow+\infty}\bm{\varepsilon}(L)$
$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{PSL}_{d}(q)$ $d\geq 2$ $d^{2}-1$ $d-1$ $\dfrac{1}{d+1}$
$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{PSU}_{d}(q)$ $d\geq 3$ $d^{2}-1$ $d-1$ $\dfrac{1}{d+1}$
$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{PSp}_{2d}(q)$ $d\geq 2$, $q$ even $2d^{2}+d$ $12d-$ 1 $\dfrac{2d-1}{2d^{2}+d}$
$d\geq 2$, $q$ odd $2d^{2}+d$ $d$ $\dfrac{1}{2d+1}$
$\dfrac{1^{1}}{1_{1_{1}}}$ $\Omega_{2d+1}(q)$ $d\geq 3$, $q$ odd $2d^{2}+d$ $22d-$ 2 $\dfrac{2d-2}{2d^{2}+d}$
$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{P}\Omega^{\pm}_{2d}(q)$ $d\geq 4$ $2d^{2}-d$ $32d-$ 3 $\dfrac{2d-3}{2d^{2}-d}$
Table 3. Asymptotic quasirandom degree of simple classical groups.
$\dfrac{1^{1}}{1_{1_{1}}}$ $L$ Comments $\sim\log_{q}|L|$ $\sim\log_{q}\mathbf{D}_{0}(L)$ $\lim\limits_{q\rightarrow+\infty}\bm{\varepsilon}(L)$
$\dfrac{1^{1}}{1_{1_{1}}}$ $E_{6}(q)$ $78$ $11$ $\dfrac{11}{78}$
$\dfrac{1^{1}}{1_{1_{1}}}$ $E_{7}(q)$ $133$ $17$ $\dfrac{17}{133}$
$\dfrac{c1^{1}}{1_{1_{1}}}$ $E_{8}(q)$ $248$ $29$ $\dfrac{29}{248}$
$\dfrac{1^{1}}{1_{1_{1}}}$ $F_{4}(q)$ $q$ even $52$ $11$ $\dfrac{11}{52}$
$\dfrac{1^{1}}{1_{1_{1}}}$ $F_{4}(q)$ $q$ odd $52$ $8$ $\dfrac{2}{13}$
@@
$\dfrac{1^{1}}{1_{1_{1}}}$ ${}^{2}E_{6}(q^{2})$ $78$ $11$ $\dfrac{11}{78}$
$\dfrac{1^{1}}{1_{1_{1}}}$ ${}^{3}D_{4}(q^{3})$ $28$ $5$ $\dfrac{5}{28}$
$\dfrac{1^{1}}{1_{1_{1}}}$ ${}^{2}F_{4}(q^{2})$ $q=\sqrt{2^{2m+1}}$ $52$ $11$ $\dfrac{11}{52}$
$\dfrac{1^{1}}{1_{1_{1}}}$ ${}^{2}B_{2}(q^{2})$ $q=\sqrt{2^{2m+1}}$ $10$ $3$ $\dfrac{3}{10}$
$\dfrac{1^{1}}{1_{1_{1}}}$ ${}^{2}G_{2}(q^{2})$ $q=\sqrt{3^{2m+1}}$ $14$ $4$ $\dfrac{2}{7}$
Table 4. Asymptotic quasirandom degree of simple exceptional groups.
@@
Acknowledgments: We thank S. Eberhard for enlightening conversations on the paper, and P. Tiep for useful informations about his articles.